OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2? y=[ ? ] x + [ ? ]

OpenStudy (anonymous):

Help fast please I gotta go soon and this is a question I am having problems with

OpenStudy (anonymous):

The equation is y = mx + b, where m is the slope, and b is the y-intercept. So where would you put the slope?

OpenStudy (anonymous):

First we plug it in the point-slope form: \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value, \(x_1\) is the x-value, and \(m\) is the slope. In this case: \(x_1 = 3\) \(y_1 = 4\) \(m = 2\) Plug them in: \(y - 4 = 2(x - 3)\) Now distribute 2 into the parenthesis, can you do that?

OpenStudy (anonymous):

@elmm29 y = [slope]x + b

OpenStudy (anonymous):

@elmm29 Multiply: \(2 \times x\) \(2 \times -3\)

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

I got y=2 and x+ -2

OpenStudy (anonymous):

@iGreen